Ag2O (s), delta Hf = -31.1kJ/mol, S=121.3J/mol , Ag (s), delta Hf =0kJ/mol, S=42.55J/mol k, O2(g) delta Hf = 0kJ/mol , S =205 j/mol K , Calculate the temperature a which delta G = 0 for the reaction Ag2O --> 2Ag + 1/2 O2 . assume that since the physical states do not change , delta H and delta S are independent of temperature between -50 C and 950C
1
Expert's answer
2019-07-01T03:55:01-0400

Solution.

"\Delta G = \Delta H - T \times \Delta S"

"\Delta G = 0"

"\Delta H = T \times \Delta S"

"T = \frac{\Delta H}{\Delta S}"

"\Delta H = (2 \times \Delta H(Ag) + 0.5 \times \Delta H(O2)) - \Delta H(Ag2O)"

"\Delta S = (2 \times \Delta S(Ag) + 0.5 \times \Delta S(O2)) - \Delta S(Ag2O)"

"\Delta H = 31.1 \frac{J}{mole}"

"\Delta S = 66.3 \frac{J}{mole \times K}"

T = 0.469 K

T = -272.681 C

Answer:

T = -272.681 C